# Human-in-the-Loop Optimization with Model-Informed Priors

### v1

@@ 4.4. Element 2: Neural Acquisition Function (naf)
The naf model in NAF+ is a neural network trained to generate acquisition values for a specific interaction scenario, learned through interactions with synthetic users. We denote the naf model as $\alpha_{t,\theta}$, where $t$ indicates the current timestep and $\theta$ corresponsd to the neural network parameters. Our training process follows the general framework introduced in the original NAF paper (Volpp et al., 2019), with key extensions tailored to the requirements of HCI tasks.
@@
The output of naf is a categorical distribution over the $s$ candidate points, denoted as $\text{CAT}[\text{nv}_{x_{1}},\text{nv}_{x_{2}},\dots,\text{nv}_{x_{s}}]$. Essentially, each $\text{nv}_{x_{i}}$ value represents the naf-generated acquisition value at a given design candidate $x_{i}$ across the full grid $x_{s}$. During training, a candidate design is sampled from this distribution to encourage exploration. During deployment, however, the design with the highest predicted acquisition probability is selected deterministically to maximize performance.
4.5. Element 3: Expected Improvement (EI)
Expected Improvement (EI) serves as the fallback acquisition strategy within NAF+. EI is one of the most widely adopted acquisition functions in Bayesian Optimization due to its simplicity, closed-form expression, and robust performance in balancing exploration and exploitation. Formally, EI at a candidate point $x$ is defined as:
(2) $\text{EI}(x)=\mathbb{E}[\max(0,f(x)-f(x^{+}))],$
where $f(x^{+})$ denotes the best observed objective value so far.
In NAF+, we apply EI to each discretized candidate $x_{i}\in\{x_{s}\}$ in the design space, producing a parallel set of acquisition scores $\text{EI}(x_{i})$. These scores are then combined with the values generated by the neural acquisition function (naf) to form the final acquisition values used for decision-making. The weighting between naf and EI is dynamically adjusted based on a novelty-aware mechanism, described in the following section, ensuring robustness when encountering out-of-distribution users.
@@ 4.6. Element 4: Novelty Detector
The novelty detector in NAF+ is a probabilistic module that estimates how likely a new user is to be out-of-distribution (i.e., significantly different from all previously seen synthetic users). Its primary role is to dynamically determine the weighting between the neural acquisition function (naf) and the fallback strategy (Expected Improvement, EI) when selecting the next design candidate.
@@ 4.7. Training and Deployment
Note that the EI and novelty detector components do not operate during the training phase. Only naf is trained by learning a policy to select promising design candidates based on inputs from the GP surrogate model. As discussed earlier, the training process is formulated as a RL problem, where the agent (i.e., naf) learns through repeated optimization episodes with synthetic users. Table 1 summarizes the RL formulation used in our meta-BO training pipeline. After training across diverse synthetic users, we obtain a large collection of design-performance pairs $(x,y)$ representing population-level behavior. These are then used to train the novelty detector (a Bayesian Neural Network, or BNN), which learns to estimate how likely a real user is out-of-distribution based on deviations from the synthetic population.
During deployment, all four components of NAF+ operate in concert. The surrogate model (GP) continuously fits the incoming data from real users. The naf model uses the updated GP predictions and the current objective weights to generate acquisition scores tailored to the user and design objectives. Simultaneously, the novelty detector assesses the novelty of the current user and dynamically adjusts the weight assigned to EI, which serves as a fallback acquisition function. The final acquisition value for each design candidate is computed as a weighted combination of naf and EI outputs, with the novelty-based weight $\lambda_{\text{EI}}$ governing the balance. We set a minimum iteration threshold $K$, so that the weighted-sum mechanism for acquisition values is activated only after $K$ iterations, ensuring that naf maintains full control during the initial phase.
RL Component NAF+ Equivalent
Episode Optimization episode with a synthetic user
Episode length $T$ Optimization budget $T$
State $s_{t}$ $\mu_{t}(x_{s}),\sigma_{t}(x_{s}),t/T,w_{1},\dots,w_{N}]$
Action $a_{t}$ Selecting a design candidate $x_{t}\in\mathcal{X}$
Reward $r_{t}$ $(f_{\text{max}}-f(x_{t}))$ (negative regret)
Transition $p(s_{t+1}|s_{t},a_{t})$ Noisy observation of $f(x_{t})$, GP update
Table 1. The RL formulation and mapping of NAF+.
@@ 4.8. Synthetic tests
Before conducting the user study, we designed two synthetic evaluations to systematically validate NAF+. These tests serve as controlled environments to probe capabilities that are difficult to isolate in live studies. Specifically, we focused on three goals: (G1) benchmarking against established baselines (standard BO and TAF), (G2) verifying that NAF+ can adapt to dynamic trade-offs between multiple objectives via explicit weight inputs, and (G3) evaluating robustness when encountering users whose behavior falls outside the training distribution. To cover both abstraction and realism, we used two tasks: a controlled double-Sphere benchmark (Test 1) for clean comparisons, and a semi-realistic typing simulation (Test 2) grounded in Fitts’ law and touch error models on touchscreen to mirror our target VR keyboard task in user study.
@@
Across both tests, NAF+ consistently demonstrated advantages over baselines. For G1 (baseline comparison), it converged faster than standard BO and TAF, particularly in the early iterations where efficiency is most critical. For G2 (dynamic objective weighting), conditioning on explicit weight inputs enabled NAF+ to adapt its search strategy more effectively than the ablated variant without weights, confirming its ability to handle shifting priorities in multi-objective optimization. For G3 (robustness to novel users), the novelty-aware fallback proved essential: when synthetic users were sampled outside the training distribution, NAF+ recovered and improved steadily, while the no-fallback variant stalled and BO/TAF required more iterations to catch up. Together, these results establish that NAF+ combines sample efficiency, flexibility, and robustness, making it well-suited for real-world HILO where the weights on multiple objective can be adjusted dynamically and users can vary dramatically.
@@ 5. User Study: NAF+ for Mid-Air Keyboard Adaptation
We conducted a user study to evaluate the effectiveness of NAF+ in a classic HCI task: mid-air typing in virtual reality. Participants completed ten iterations of typing (each consisting of one to two sentences) with adaptive keyboards generated by three methods. Subsections 5.1–5.4 describe the study design, including the experimental conditions (methods), the user models used to train NAF+, the objective function, and the study apparatus. The study was conducted in two phases. Phase 1 (Subsection 5.5) provided data to pre-train two optimization methods, TAF and NAF+. Subsection 5.6 details the implementation and training procedure of NAF+. Phase 2, the final evaluation, is described in Subsections 5.7 (study setup) and 5.8 (results). Finally, Subsection 5.9 presents our discussion of the study findings.
@@ 5.2. Typing Simulation for Training NAF+
To pre-train NAF+, we implemented a generative simulation of sentence typing that models the motor execution and touch uncertainty of a human typist. The simulator combines two established models of touchscreen interaction: Fitts’ Law for movement time estimation (MacKenzie and Buxton, 1992) and the Dual Gaussian Distribution (DGD) model for touch accuracy (Bi et al., 2013; Bi and Zhai, 2016, 2013).
@@ 5.3. Objective Function and Task Setting
The simulator outputs two primary metrics: typing speed (words per minute, WPM) and error rate. To combine them into a single scalar measure, we normalize each into unitless functions of Speed and Accuracy.
@@ 5.4. Study Apparatus and Implementation
The virtual keyboard and typing experiment were implemented in Unity and executed on a Windows 10 desktop with an NVIDIA GeForce RTX 3050 GPU. We built the keyboard on top of the MRTK keyboard framework [^2] https://github.com/microsoft/MixedRealityToolkit-Unity, which we customized to support adaptive resizing of key height and width. The study was conducted on a Meta Quest Pro headset. Target sentences were sampled from the Enron Mobile Email Dataset, with special characters removed and multiple sentences concatenated to ensure a fixed length of 26 characters (excluding the space key) per iteration.
@@ 5.6. Implementing and Training NAF+
Before Phase 2 of the user study, we prepared NAF+ by generating synthetic users, implementing its components, and training the model.
@@ 5.7.1. Study Design and Procedure
The study employed a within-subject design with 12 participants (5 female, age 24–33). Each participant completed three blocks, one per optimization method, with order counterbalanced using a Latin square. In each block, participants typed 10 sentences (iterations), following a short practice session of 10 iterations. Between blocks, a 5-minute break was provided to reduce fatigue. The entire session lasted less than 60 minutes per participant. Finally, the participants were asked to fill in the raw NASA-TLX questionnaire.
@@ 5.7.2. Methods
We described the implementation of NAF+ in Section 5.6. The novelty detector automatically selected the correct objective weight combination ($w_{\text{speed}}=0.7$, $w_{\text{accuracy}}=0.3$).
Transfer Acquisition Function (TAF).
@@ 5.7.3. Results
Figure 3. The weighted-sum performance achieved by different conditions (optimization methods) at each iteration. This plot presents the running best performance (i.e., the best performance of each participant achieved up to this iteration), which is a common way of analyzing iterative optimization tasks. The * sign indicates significant differences ($p<0.05$). The thicker, black line in each box shows the mean value, and the thin, orange line indicates the median value.
Figure 4. The typing speed at each iteration, which is then contributed to the calculation of the final objective function as described in subsection 5.3. The thicker, black line in each box shows the mean value, and the thin, orange line indicates the median value.
Figure 5. The error rate at each iteration, which is then contributed to the calculation of the final objective function as described in subsection 5.3. The thicker, black line in each box shows the mean value, and the thin, orange line indicates the median value.
@@ 5.8. Findings and Discussion
In this study, we investigated three optimization methods for HILO for mid-air keyboard adaptation. Particularly, we aim to understand how different strategies for leveraging prior knowledge influence performance over successive iterations. The key findings are that all three methods led to performance improvements across iterations, yet the resulting performances diverged in the early iterations: NAF+ provided an advantage in the second and third iterations.
@@
TAF’s performance is significantly poorer in early iterations, which can be attributed to the limited coverage of its prior models. The five priors collected in Phase 1 could not sufficiently capture the diversity of population performance, leading to poor prediction in unexplored regions of the design space. In principle, expanding the number of priors could improve coverage and even approach optimal results. However, there are two fundamental challenges. First, there is no clear principle for predicting how many prior users are sufficient, which makes it difficult to guarantee generalization. Second, scaling up the number of priors linearly increases computation during deployment, rendering this approach less practical in real-world scenarios. By contrast, NAF+ has the unique strength of not being limited to a finite set of prior users. Instead, it can generate unlimited synthetic users by sampling from learned parameter distributions. This synthetic population enables extensive offline training before deployment, resulting in more consistent early-iteration performance. In addition, the novelty detector allows NAF+ to flexibly adjust the weighting between the NAF policy and the EI values calculated from the current user’s model. This mechanism provides robustness in out-of-distribution cases: when priors are less informative, NAF+ behaves similarly to TAF by relying more heavily on the current user’s data, ensuring at least comparable performance.
@@
ConBO showed promising overall performance but was less effective in the earliest iterations, primarily because the first few participants exhibited relatively poor performance during their initial trials. The fundamental challenge of ConBO is that it learns gradually across users. For the first few participants, the population model contains little information, so its search behavior resembles random exploration. Once enough participants are aggregated, however, the model improves steadily. This trend is evident when comparing the performance of early and later groups of participants: Group 1 (P1–3) = 0.37 (std = 0.12), Group 2 = 0.61 (std = 0.07), Group 3 = 0.59 (std = 0.19), and Group 4 = 0.63 (std = 0.04). These results clearly show that ConBO benefits from accumulating user data, but at the expense of initial participants who experience suboptimal adaptation. In contrast, NAF+ does not suffer from this early-stage drawback, as its model is fully trained prior to deployment using synthetic users. The group-wise performances further illustrate this stability: Group 1 = 0.54 (std = 0.10), Group 2 = 0.56 (std = 0.12), Group 3 = 0.66 (std = 0.25), and Group 4 = 0.65 (std = 0.03). Thus, NAF+ provides stronger early performance and avoids the cold-start problem observed in ConBO.
@@
In conclusion, this study highlights the trade-offs between different strategies for leveraging prior knowledge in HILO. TAF is limited by finite priors and scalability concerns, ConBO improves gradually but suffers from early-user effectiveness, and NAF+ combines the advantages of prior-based learning with scalable synthetic-user generation to achieve stronger early adaptation. These findings suggest that simulation-driven approaches such as NAF+ offer a promising path toward scalable, generalizable, and efficient human-in-the-loop optimization for adaptive interfaces.

### v2

@@ 4.4. Element 2: Neural Acquisition Function (naf)
The naf model in NAF+ is a neural network trained to generate acquisition values for a specific interaction scenario, learned through interactions with synthetic users. We denote the naf model as $\alpha_{t,\theta}$, where $t$ indicates the current timestep and $\theta$ corresponds to the neural network parameters. Our training process follows the general framework introduced in the original NAF paper (Volpp et al., 2019), with key extensions tailored to the requirements of HCI tasks.
@@
The output of naf is a categorical distribution over the $s$ candidate points, denoted as $\text{CAT}[\text{nv}_{x_{1}},\text{nv}_{x_{2}},\dots,\text{nv}_{x_{s}}]$. Essentially, each $\text{nv}_{x_{i}}$ value represents the naf-generated acquisition value at a given design candidate $x_{i}$ across the full grid $x_{s}$. Table 1 summarizes the RL formulation used in our meta-BO training pipeline. During training, a candidate design is sampled from this distribution to encourage exploration. During deployment, however, the design with the highest predicted acquisition probability is selected deterministically to maximize performance.
RL Component NAF+ Equivalent
Episode Optimization episode with a synthetic user
Episode length $T$ Optimization budget $T$
State $s_{t}$ $\mu_{t}(x_{s}),\sigma_{t}(x_{s}),t/T,w_{1},\dots,w_{N}]$
Action $a_{t}$ Selecting a design candidate $x_{t}\in\mathcal{X}$
Reward $r_{t}$ $(f_{\text{max}}-f(x_{t}))$ (negative regret)
Transition $p(s_{t+1}|s_{t},a_{t})$ Noisy observation of $f(x_{t})$, GP update
Table 1. The RL formulation and mapping of NAF+.
4.5. Element 3: Expected Improvement
A typical EI serves as the fallback acquisition strategy, with details provided in subsection 4.1. We apply EI to each discretized candidate $x_{i}\in\{x_{s}\}$ in the design space, producing a parallel set of acquisition scores $\text{EI}(x_{i})$. These scores are then combined with the values generated by the neural acquisition function (naf) to form the final acquisition values used for decision-making. The weighting between naf and EI is dynamically adjusted based on a novelty-aware mechanism, described in the following section, ensuring robustness when encountering out-of-distribution users.
@@ 4.6. Element 4: Novelty Detector
The novelty detector in NAF+ is a probabilistic module that estimates how likely a new user is to be out-of-distribution (i.e., significantly different from all previously seen synthetic users). Its primary role is to dynamically determine the weighting between the neural acquisition function (naf) and the fallback strategy (EI) when selecting the next design candidate.
@@ 4.7. Training and Deployment
naf is trained with a large number of synthetic users. Then, we obtain a large collection of design-performance pairs $(x,y)$ representing population-level behavior, which are used to train the novelty detector. During deployment, the novelty detector assesses the novelty of the user and adjusts the weight assigned to EI, as described in Equation 4. We set a minimum iteration threshold $K=3$, so that the weighted-sum mechanism for acquisition values is activated from the 4th iterations, ensuring that naf maintains full control during the initial 3 iterations.
@@ 4.8. Synthetic tests
Before conducting the user study, we designed two synthetic evaluations to systematically validate NAF+. These tests serve as controlled environments to probe capabilities that are difficult to isolate in live studies. Specifically, we focused on three goals: (G1) benchmarking against established baselines: standard BO and Transfer Acquisition Function, which is an established meta-BO method used in prior HCI works (Wistuba et al., 2018; Liao et al., 2024a); (G2) verifying that NAF+ can adapt to dynamic trade-offs between multiple objectives via explicit weight inputs; (G3) evaluating robustness when encountering users whose behavior falls outside the training distribution. To cover both abstraction and realism, we used two tasks: a controlled double-Sphere benchmark (Test 1) for clean comparisons, and a semi-realistic typing simulation (Test 2) grounded in Fitts’ law and touch error models on a touchscreen to mirror our target VR keyboard task in the user study.
@@
Across both tests, NAF+ consistently demonstrated advantages over baselines. For G1 (baseline comparison), it converged faster than standard BO and TAF, particularly in the early iterations where efficiency is most critical. For G2 (dynamic objective weighting), conditioning on explicit weight inputs enabled NAF+ to adapt its search strategy more effectively than the ablated variant without weights, confirming its ability to handle shifting priorities in multi-objective optimization. For G3 (robustness to novel users), the novelty-aware fallback proved essential: when synthetic users were sampled outside the training distribution, NAF+ recovered and improved steadily, while the no-fallback variant stalled and BO/TAF required more iterations to catch up. Together, these results establish that NAF+ combines sample efficiency, flexibility, and robustness, making it well-suited for real-world HILO where the weights on multiple objective can be adjusted dynamically and users can vary dramatically. The results of the synthetic tests are presented in the Appendix.
@@ 5. User Study: NAF+ for Mid-Air Keyboard Adaptation
We conducted a user study to evaluate the effectiveness of NAF+ in a classic HCI task: mid-air typing in virtual reality; the interaction is shown in Figure 3. Participants completed ten iterations of typing (each consisting of one to two sentences) with adaptive keyboards with varying key height and width, generated by three methods. Subsections 5.1–5.4 describe the study design, including the experimental conditions (methods), the user models used to train NAF+, the objective function, and the study apparatus. The study was conducted in two phases. Phase 1 (Subsection 5.5) provided data to pre-train two optimization methods, TAF and NAF+. Subsection 5.6 details the implementation and training procedure of NAF+. Phase 2, the final evaluation, is described in Subsections 5.7 (study setup) and 5.8 (results). Finally, Subsection 5.9 presents our discussion of the study findings.
@@ 5.2. Typing Simulation for Training NAF+
To train NAF+, we implemented a generative simulation of sentence typing that models the motor execution and touch uncertainty of a human typist. The simulator combines two established models of touchscreen interaction: Fitts’ Law for movement time estimation (MacKenzie and Buxton, 1992) and the Dual Gaussian Distribution (DGD) model for touch accuracy (Bi et al., 2013; Bi and Zhai, 2016, 2013).
@@
Figure 3. Our user study aims to personalize the mid-air keyboard’s key height and width using three methods: our NAF+, Transfer Acquisition Function (TAF), and Continual Bayesian optimization (ConBO). TAF and ConBO are both existing BO approaches for achieving efficient HILO with prior experience. Here we present some possible keyboard variants within the design parameter range.
@@ 5.3. Objective Function and Task Setting
Our goal is to optimize the key’s height and width of mid-air keyboards. The simulator outputs two primary metrics: typing speed (words per minute, WPM) and error rate. To combine them into a single scalar measure, we normalize each into unitless functions of Speed and Accuracy.
@@ 5.4. Study Apparatus and Implementation
The virtual keyboard and typing experiment were implemented in Unity and executed on a Windows 10 desktop with an NVIDIA GeForce RTX 3050 GPU. We built the keyboard on top of the MRTK keyboard framework [^2] https://github.com/microsoft/MixedRealityToolkit-Unity, which we customized to support adaptive resizing of key height and width. The study was conducted on a Meta Quest Pro headset, as shown in Figure 3. Target sentences were sampled from the Enron Mobile Email Dataset, with special characters removed and multiple sentences concatenated to ensure a fixed length of 26 characters (excluding the space key) per iteration.
@@ 5.6. Implementing and Training NAF+
Before Phase 2 of the user study, we prepared NAF+ by generating synthetic users, and implementing and training its components.
@@ 5.7.1. Study Design and Procedure
The study employed a within-subject design with 12 participants (5 female, age 24–33). Each participant completed three blocks, one per optimization method, with order counterbalanced using a Latin square. In each block, participants typed 10 sentences (iterations). Before any block, the participants had a short practice session of 10 sentences where the keyboard dimension changed randomly. Between blocks, a 5-minute break was provided to reduce fatigue. The entire session lasted less than 60 minutes per participant. Finally, the participants were asked to fill in the raw NASA-TLX questionnaire followed by an open-ended interview.
@@ 5.7.2. Methods
We described the implementation of NAF+ in Section 5.6. The novelty detector automatically selected the correct objective weight combination ($w_{\text{speed}}=0.7$, $w_{\text{accuracy}}=0.3$). We set the novelty detector threshold $\tau=0.1$. This choice aligns with common statistical practice, where a p-value below $0.1$ is often interpreted as an indicator of outliers or model mismatch. Because our novelty estimate is based on the running average of p-values across iterations, a slightly more permissive threshold also ensures that the system can detect meaningful deviations early. This is also the value we used in the synthetic tests, shown to be an effective setting. Importantly, the novelty detector begins at the fourth iteration. In the initial three iterations, EI weights are set to be 0, and the weights assigned on naf is set to be $1$.
Transfer Acquisition Function (TAF)
@@ 5.7.3. Results
Figure 5 and Figure 6 show the individual performance metrics throughout the iterations achieved by three conditions. We ran one-way repeated-measures ANOVA on each iteration across all metrics and found that the NAF+ generally led to either comparable or significantly better performances than the other baselines. We denote the significant differences in the figures.
@@
Figure 4. The weighted-sum performance achieved by different conditions (optimization methods) at each iteration. This plot presents the running best performance (i.e., the best performance of each participant achieved up to this iteration), which is a common way of analyzing iterative optimization tasks. The * sign indicates significant differences ($p<0.05$). The thicker, black line in each box shows the mean value, and the thin, orange line indicates the median value. The results showed that our NAF+ enables faster convergence than other approaches. In particular, NAF+ achieved statistically better performance than TAF in the second and third iterations, and outperformed ConBO in the third iteration.
Figure 5. The typing speed at each iteration, which is then contributed to the calculation of the final objective function as described in subsection 5.3. The * sign indicates significant differences ($p<0.05$). The thicker, black line in each box shows the mean value, and the thin, orange line indicates the median value. The results showed that NAF+ enabled faster typing speed than TAF in the second and third iterations, and faster than ConBO in the third iteration.
Figure 6. The error rate at each iteration, which is then contributed to the calculation of the final objective function as described in subsection 5.3. The * sign indicates significant differences ($p<0.05$). The thicker, black line in each box shows the mean value, and the thin, orange line indicates the median value. The results showed that NAF+ led to lower error rates than the other methods in the third iteration.
@@ 5.8. Findings and Discussion
5.8.1. Findings on Performance
@@
Three methods led to performance improvements across iterations, yet the resulting performances diverged in the early iterations: NAF+ provided an advantage in the second and third iterations. TAF’s performance is significantly poorer in early iterations, which can be attributed to the limited coverage of its prior models. The five priors collected in Phase 1 could not sufficiently capture the diversity of population performance, leading to poor prediction in unexplored regions of the design space. In principle, expanding the number of priors could improve coverage and even approach optimal results. However, there are two fundamental challenges. First, there is no clear principle for predicting how many prior users are sufficient, which makes it difficult to guarantee generalization. Second, scaling up the number of priors linearly increases computation during deployment, rendering this approach less practical in real-world scenarios. By contrast, NAF+ has the unique strength of not being limited to a finite set of prior users. Instead, it can generate unlimited synthetic users by sampling from learned parameter distributions. This synthetic population enables extensive offline training before deployment, resulting in more consistent early-iteration performance. In addition, the novelty detector allows NAF+ to flexibly adjust the weighting between the NAF policy and the EI values calculated from the current user’s model. This mechanism provides robustness in out-of-distribution cases: when priors are less informative, NAF+ behaves similarly to TAF by relying more heavily on the current user’s data, ensuring at least comparable performance.
@@
ConBO showed promising overall performance but was less effective in the earliest iterations, primarily because the first few participants exhibited relatively poor performance during their initial trials. The fundamental challenge of ConBO is that it learns gradually across users. For the first few participants, the population model contains little information, so its search behavior resembles random exploration. Once enough participants are aggregated, however, the model improves steadily. This trend is evident when comparing the performance of early and later groups of participants: Group 1 (P1–3) = 0.37 (std = 0.12), Group 2 = 0.61 (std = 0.07), Group 3 = 0.59 (std = 0.19), and Group 4 = 0.63 (std = 0.04). These results clearly show that ConBO benefits from accumulating user data, but at the expense of initial participants who experience suboptimal adaptation. In contrast, NAF+ does not suffer from this early-stage drawback, as its model is fully trained prior to deployment using synthetic users. The group-wise performances further illustrate this stability: Group 1 = 0.54 (std = 0.10), Group 2 = 0.56 (std = 0.12), Group 3 = 0.66 (std = 0.25), and Group 4 = 0.65 (std = 0.03). Thus, NAF+ provides better performance for the initial participants compared to ConBO.
@@
To summarize, TAF is limited by finite priors and scalability concerns, ConBO improves gradually but suffers from early-user effectiveness, and NAF+ combines the advantages of prior-based learning with scalable synthetic-user generation to achieve stronger early adaptation.
5.8.2. Importance of the novelty detector and EI
We further analyzed the contribution of EI and novelty detector by examining the EI weights of all the participants across iterations 4-10 (at which weighted-sum mechanism performs). The EI weights are shown in Figure 7. Overall, we observe a gradual upward trend in EI weights, indicating that as the optimization progresses, the optimization increasingly relies on the EI component. A closer examination of individual EI-weight trajectories reveals different participant groups that have distinct levels of novelty in performance characteristics: A few participants well-modeled users, who consistently show low and stable EI weights (e.g., 0.3–0.5). These users behave very similarly to the predictive model’s expected performance distributions. A few participants are novel users. They show consistently high EI weights, often rising rapidly toward the upper range (e.g., 0.7–1.0). Their performance profiles deviate substantially from what the predictive model expects. Yet, most participants have moderate EI weights with small oscillations, indicating most participants partly match the model’s expectations, which is a natural outcome.
These observations show the importance of the novelty detector and the EI weighting mechanism. With them, NAF+ can effectively dynamically adjusts how strongly it should rely on learned priors versus current observations, maintaining efficiency without sacrificing personalization across diverse users.
Figure 7. The evolution of EI weights ($\lambda_{EI}$) using NAF+ at each iteration. The thicker, black line in each box shows the mean value, and the thin, orange line indicates the median value. Overall, the novelty levels increase across iterations, and participants demonstrate high differences from very novel users (behaviors largely differ from predictions) to well-modeled users (behaviors are highly aligned with predictions). This shows the importance of employing the novelty detector and EI as a fallback acquisition.
5.8.3. Qualitative analysis of the user experience with ConBO
While no significant differences were found in NASA-TLX questions, the first few participants in ConBO shared more frustrated experiences. As previously analyzed, Group 1 (P1–P3) in ConBO exhibited worse overall performance compared to later groups, indicating that ConBO was still accumulating knowledge and had not yet formed a reliable prior for personalization. In the interviews, Group 1 participants described the initial keyboards as feeling “weird,” with “impossible ratios” or being “way too small for me to type comfortably.” They also expressed that the system did not seem to adapt meaningfully to their behavior. For example, P2 said, “I felt the keyboard randomly changes dimensions,” and P1 noted, “It did not improve from where I did better.” In contrast, participants in later groups described a smoother experience. Notably, complaints about impossible layouts or misaligned adaptations were no longer observed, nor in the other conditions. Overall, these observations confirm a clear drawback of ConBO: its performance for early participants is less reliable, as the system is still accumulating experience and has not yet formed a prior for effective optimization.